OpenStudy (anonymous):

(1/5)+(7/5)y=x 10x-14y=2 How do I make it in slope intercept form?

OpenStudy (gorv):

slope intercept form is y=mx+c so keep y alone on one side and move others on other side

OpenStudy (gorv):

@Kamilchehayeb

OpenStudy (gorv):

(7/5)y=x-1/5

OpenStudy (gorv):

y=5/7x-1/7

OpenStudy (anonymous):

I know, but I need the correct answer, I got different ones.

OpenStudy (gorv):

like this do 2nd

OpenStudy (anonymous):

-10x+(1/7)?

OpenStudy (anonymous):

gorv got the 1st one right here's the second 10x-14y=2 minus 10x -14y=-10x+2 divide by -14 y=(5/7)x-(1/7)

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

they are the same line

OpenStudy (gorv):

10 will also divided by 14 like 2

OpenStudy (anonymous):

Oh ok, thanks guy @mathman806 ans @gorv

OpenStudy (anonymous):

no prob